Please help me solve this equation..please write without exponents 16^3/4
Answer by stanbon(75887)

Please help me solve this equation..please write without exponents 16^3/4
------------------
That exponent tells you to take the 4th root and to cube the result.
--------------------------------------
So follow the instructions:
16^(3/4) = 2^3 = 8
============================
Cheers,
Stan H.